# FIDLAR: Forecast-Informed Deep Learning Architecture for Flood Mitigation

### v1

@@ 5 Experiments
5.1 Data
We obtained data from the South Florida Water Management District’s (SFWMD) DBHydro database District (2023) for the coastal stretch of the South Florida watershed. The data set consists of hourly observations for water levels and external covariates from Jan. 1st, 2010 to Dec. 31, 2020, a period totaling eleven years. As shown in Figure 6, the river system under consideration has two branches and includes several hydraulic structures (gates, pumps) to control water flows. Water levels in coastal systems are also impacted by tides. We aim to predict effective schedules on hydraulic structures (gates, pumps) to minimize flood risks at four specific locations marked by green circles in Figure 6. To highlight the risk posed by floods, we note that this portion of the river system flows through Miami downtown, with a sizable population, many commercial enterprises, and an international airport in its vicinity.
Figure 6: Schemetic diagram of study domain - South Florida. There are three water stations with hydraulic structures (S26, S25B, S25A), one simple water station, S1 (green circle in the middle), and one station monitoring tide information from the ocean.
Table 1: Summary of the data set (Jan. 1, 2010 $\sim$ Dec. 31, 2020).
@@ 5.1 Data
5.2 Experimental designs
The sliding input window Li et al. (2014) (also known as look-back window Gidea and Katz (2018) strategy was used to process the entire dataset Shi et al. (2023a)), which helps to ensure that the input and output vector lengths are constant. The datasets were split in chronological order with the first 80% for training and the remaining 20% for testing.
@@ 5.2 Experimental designs
The primary role of the Flood Evaluator is to forecast flood events by predicting water levels for given input conditions. To assess its performance, we compared several DL models for the evaluation, including our proposed GTN model as well as a physics-based model (HEC-RAS). We measured accuracy using multiple metrics: (a) the mean absolute error (MAE), (b) the root mean square error (RMSE) computed between the predicted and actual water levels, (c) the number of time points where the upper or lower thresholds were breached, and (d) the area between the water level curve and the threshold bars. For consistency, we used a look-back window of length $w=72$ hours and a prediction window of length $k=24$ hours. The eight DL methods used are listed below (more details in Appendix B):
  • MLP Suykens et al. (1995): Multilayer perceptron can learn complex non-linear dependencies;
  • RNN Medsker and Jain (2001): Recurrent neural networks are good at processing sequential data;
  • CNN O’Shea and Nash (2015): A 1D convolutional neural network;
  • GNN Kipf and Welling (2016): Graph neural network with nodes representing variables and edges representing spatial dependencies;
  • TCN Bai et al. (2018): Temporal dilated convolutional network with an exponentially large receptive field;
  • RCNN Zhang and Dong (2020): Combined RNN and CNN model for time series forecasting;
  • Transformer Vaswani et al. (2017): Attention-based network for sequence modeling (only encoder);
  • GTN: Combining GNNs with LSTMs, CNNs, and transformers, as described in Figure 5.
Flood Mitigation with FIDLAr.
FIDLAr requires both Evaluator and Manager components and we chose the GTN model for the Evaluator, since it was the best performing model. For the Manager model, we experimented with one rule-based method, and two genetic algorithms – one with a physics-based HEC-RAS evaluator Leon et al. (2020) and one with our DL-based GTN evaluator, and several DL-based managers. As shown in Table 3, the list of DL approaches used in our experiments for the Manager included MLP, RNN, CNN, GNN, TCN, RCNN, Transformer, and GTN. Performance of FIDLAr was measured using (a) the number of time steps where the upper/lower thresholds were exceeded for the water levels, and (b) the area between the water level curve and the threshold bars.
5.3 Analysis of computational time
Since FIDLAr was designed for real-time flood control, we measured the running times of the models used in this work.
Table 2: Comparison of model performances for the Flood Evaluator on the test set, specifically at time t+1 for measurement station S1. The terms “Over Timesteps” and “Under Timesteps” indicate the number of time steps during which water levels exceed the upper threshold or fall below the lower threshold, respectively. Similarly, “Over Area” and “Under Area” pertain to the area between the water level curve and upper or lower threshold, as was illustrated in Figure 4. Results in red are lowest in that column or closest to the ground truth (in blue).
@@ 5.3 Analysis of computational time
Table 3: Comparison of model performances for the Flood Manager on the test set, specifically at time t+1 for measurement station S1. The $*{}^{*}$ denotes that the GA method was used with a physics-based (HEC-RAS) evaluator. The -- denotes that the experiments were timed out. The ${}^{{\dagger}}$ denotes that the GA method was used with the GTN as the evaluator. All other rows are DL-based flood managers with a DL-based GTN as the evaluator. Results in red font are the best in that column.
@@
GA-based Genetic $*{}^{*}$ (GA) - - - -
Genetic ${}^{{\dagger}}$ (GA) 86 16.54 454 104
@@
6 Experimental Results
6.1 Flood Prediction
To evaluate the accuracy of the Flood Evaluator, we test FIDLAr with various DL models on the test set by predicting the water levels $k=24$ hours in advance. Table 2 shows the MAEs and RMSEs for the DL-based models and the physics-based model (HEC-RAS). As discussed earlier, we also analyze the time steps and areas over and under the thresholds. We set the upper threshold (flood level) at 3.5 feet and the lower threshold (wastage level) at 0.0 feet. However, the methods remain consistent for many reasonable choices of threshold values. From Table 2, we note that GTN surpasses the performance of other models with predictions most closely aligned with the ground truth (highlighted in red). More results are provided in Table 8 in Appendix C.1.
6.2 Flood Mitigation
Performance measures.
Table 3 shows that all DL-based methods consistently performed better for site S1 than rule-based and GA-based approaches. Moreover, GTN has the best performance under all four metrics, whether it is to control floods or water wastage. The results for all other measurement sites follow a similar pattern, as shown in Table 9 in Appendix C.2.
Performance visualization.
We visualize the mitigated water levels for a short sample spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest. Figure 7 indicates that FIDLAr equipped with GTN has resulted in water levels within the upper and lower thresholds. The zoomed portion shows a 2.5-hour period where the water levels are decreased under 3.5 feet based on the predicted gate and pump schedules. We provide the corresponding performance in Table 10. More visuals are in Appendix D.
Figure 7: Visualization of water levels with various methods for flood mitigation. We zoomed in the period of $t=2\sim 4.5$ in gray.
@@ 6.3 Ablation study
The ablation study quantifies the contribution of each of the components of GTN, by measuring the performance of GTN after removing each of the individual components. Results are shown in Table 4.
Table 4: Ablation study for flood mitigation for the entire test set (for time point t+1 at S1). The last row indicates the performance of the FIDLAr system with GTN as proposed in Figure 5. The best results are in red font.
Removed Over Over Under Under
@@
CNN 37 4.37 476 85.54
Transformer 32 3.57 325 57.42
GNN 56 5.90 479 86.22
LSTM 35 4.34 329 56.74
Attention 32 3.59 341 60.48
GTN 22 2.23 299 53.34
6.4 Computational time
Table 5 shows the running times for all the methods for the evaluator component and for the whole flood mitigation system in its test phase. All the DL-based approaches are several orders of magnitude faster than the currently used physics-based and GA-based approaches. The table also shows the training times for the DL-based approaches, although they are not necessary for real-time performance.
Table 5: The running time for flood prediction and mitigation. The running time for the rule-based method is not reported since historical data was directly used. $*{}^{*}$, which combines a GA-based tool and HEC-RAS, took too long and was not reported although the time for a small data set is reported in Tables 10, 11 in Appendix C.3. ${}^{{\dagger}}$, which combines the GA-based tool with GTN, also took too long but was estimated using a smaller sample.
Model Prediction Mitigation
@@ 6.4 Computational time
GTN 119 min 2.95 s 35 min 4.90 s
6.5 Model explainability
Attention-based methods allow us to calculate the “attention scores” assigned to an input variable to compute a specific output variable. This is exemplified in the heatmap shown in Figure 8, which shows the attention scores assigned to the tide (columns) to compute the gate schedule output (rows) for 24 hours into the future. Note that there are 96 columns and 24 rows because we use 72 hours of past tidal observations and 24 hours of future predicted tidal data to predict 24 hours of the gate schedule into the future. Thus the rows $[0,23]$ correspond to the 24 hours into the future, while the columns $[0,95]$ also include 72 hours of the recent past. Thus, $t=72$ corresponds to the “current” time point, and the columns $[72,95]$ correspond to the same time points as rows $[0,23]$.
Figure 8: Attention scores between tide input and one of the gate schedule outputs. $x$ and $y$ axes are the tide and gate over time.
@@ 7 Discussion
The explainability feature, which was shown with an example in Figure 8, can provide significant insights into our results. Firstly, the simple example with the attention paid to the variable “tide” shows that the brightest patches in the heatmap occur in the last 24 columns. This highlights the importance of the novel idea of using tidal forecasting information in our flood mitigation approach. Given that tidal information within a 24-hour window can be quite accurate, this is useful. While tides may have a more predictable pattern over time, the contribution of rain on the water levels can also be seen for other time points. A second critical insight is that the brightest attention patches are in columns where the tide is at its highest is critical to the prediction of gate schedules. Additionally, the water level at the first high tide peak after the “current” time is more significant than the other two. Third, the gate schedule has peaks at times $t=11h$ and $t=22h$ into the future, which correspond to the lowest points of the tide. This implies that the optimal time for pre-releasing water is during low tide phases. Opening gates during high tide periods in coastal river systems is less advisable, as it may lead to water flowing back upstream from the ocean. Finally, we observe that while there is a light patch around column $t=65h$, suggesting mild attention for the previous high tide peak, but almost no attention to any of the peaks prior to that. This again suggests that we could have chosen to use a smaller window for the past input. Doing this analysis for more time points could provide evidence for the right value of $w$, the size of the look-back window.
@@ 8 Conclusions
FIDLAr is a DL-based tool to compute water “pre-release” schedules for hydraulic structures in a river system to achieve effective and efficient flood mitigation, while ensuring that water wastage is avoided. This was made possible by the use of well-crafted loss functions for the DL models. The dual component design (with a Manager and an Evaluator) is a strength of FIDLAr. During training, the gradient-based back-propagation from the Evaluator helps to fine-tune the Manager. The reliability of forecasts of variables like precipitation and tides is well known, but their critical importance may have been best demonstrated for the first time using the attention mechanism employed herein.
All the DL-based versions of FIDLAr are several orders of magnitude faster than the (physics-based or GA-based) competitors while achieving improvement over other methods in flood mitigation. These characteristics allow us to entertain the possibility of real-time flood management, which was previously unthinkable.
@@ Acknowledgments
This work is part of the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) project, which is funded by the National Science Foundation under award number 2118329.
@@ Appendix A Data and Processing
(a) Flood Evaluator
(b) Flood Manager
Figure 9: Flood Evaluator and Flood Manager in FIDLAR.
@@ B.1 Architecture
  • GNN: We consider each time series as one node and the connection as the edge. For the variables in Fig. 6, we assign “1” to the edge between two connected nodes, otherwise, “0”. For the architecture, we stacked two graph convolutional networks (GCNs) with $32$ and $16$ neurons. In parallel, a LSTM layer with $32$ units is used to learn the temporal dynamics from the input. We then concatenate the embedding from GCN and LSTM and output the outcomes.
@@
  • Transformer: The standard Transformer encoder Vaswani et al. [2017] is used. In our case, we only used 1 Encoder with 1 head. We set head size=128, embedding dimension=64, Dense layer with 32 neurons. Dropout is used after each layer above.
  • GTN: We combine the components of GNNs, LSTMs, Transformer, CNNs, and attention, as described in Fig. 5. We have $32$ and $16$ for GNN’s channels, $16$ for LSTM units, $96$ for CNN filters, and one Transformer encoder with $3$ heads (head size=192, embedding dimension=96).
@@
When we used the above models for Flood Evaluator and Flood Manager, the input and output need to be changed accordingly (see Figures 2 and 3). To align with real-world scenarios, we enforce stringent constraints that restrict the output of Manager (values for gate and pump openings) to fall within the range of $[MIN,MAX]$ by a modified RELU activation function in Eq. (7). Fig. 10 provides the corresponding visualization.
@@ B.2 Training hyper-parameters
Table 6: Best hyperparameters of each DL model for Flood Evaluator.
@@
Table 7: Best hyperparameters of each DL model for Flood Manager.
@@ C.1 Flood prediction at all locations
In Table 2, we provided the comparison of the performance of different models for the Flood Evaluator at time point $t+1$, but at one location (S1). In Table 8, we provide the comparison of the performance of different models for the Flood Evaluator on the test set at time $t+1$, but for all locations.
@@
Table 8: Comparison of different models for the Flood Evaluator on the test set (at time t+1 for all stations). “Over Timesteps” (and “Under Timesteps”) represent the number of time steps during which the water levels exceed the upper threshold (subceed the lower threshold, resp.). Similarly, “Over Area” (and “Under Area”) refer to the area between the water level curve and the upper threshold (lower threshold, resp.), as shown in Fig. 4 in the manuscript.

### v2

@@ 6 Experiments
6.1 Dataset
We obtained data from the South Florida Water Management District’s (SFWMD) DBHydro database (District 2023) for the coastal stretch in South Florida. The data set consists of hourly observations for water levels and external covariates from January 1, 2010 to December 31, 2020. As shown in Figure 5, the river system has two branches and includes several hydraulic structures (gates, pumps) to control water flows. We aim to predict effective control schedules on hydraulic structures (gates, pumps) to minimize flood risks at four specific locations marked by green circles.
Figure 5: Schemetic diagram of study domain. There are three water stations with hydraulic structures (S26, S25B, S25A), one simple water station, S1 (green circle in the middle), and a station monitoring the tide level from the ocean.
@@ 6.1 Dataset
Table 1: Summary of the data set.
Figure 6: Graph Transformer Network for Flood Evaluator. Input variables include Rainfall, Pump, Gate, Tide, and Water levels as shown in Figure 5 and Table 1, which are generically denoted by $R,T,G,P$, and $W$. Output is water levels.
6.2 Experimental Design
The sliding input window (Li et al. 2014) (also known as look-back window (Gidea and Katz 2018) strategy was used to process the entire dataset (Shi et al. 2023a)). For consistency, we used a look-back window of length $w=72$ hours and a prediction window of length $k=24$ hours. The dataset was split in chronological order with the first 80% for training and the remaining 20% for testing. The eight DL methods below are used for Flood Manager and Flood Evaluator. We run all experiments on one NVIDIA A100 GPU with 80GB memory.
  • MLP (Suykens, Vandewalle, and De Moor 1995): Multilayer perceptron can learn non-linear dependencies;
  • RNN (Medsker and Jain 2001): Recurrent neural networks are good at processing sequential data;
  • CNN (O’Shea and Nash 2015): A 1D convolutional neural network;
  • GNN (Kipf and Welling 2016): Graph neural network with nodes representing variables and edges representing spatial dependencies;
  • TCN (Bai, Kolter, and Koltun 2018): Temporal dilated convolutional network with an exponentially large receptive field;
  • RCNN (Zhang and Dong 2020): Combined RNN and CNN model for time series forecasting;
  • Transformer (Vaswani et al. 2017): Attention-based network for sequence modeling (only encoder);
  • GTN (Ours): Combining GNNs with LSTMs, CNNs, and transformers, as described in Figure 6.
@@ 6.2 Experimental Design
The role of the Flood Evaluator is to forecast flood events by predicting water levels for given input conditions. We set the upper threshold (flood level) at 3.5 feet and the lower threshold (wastage level) at 0.0 feet. However, the methods remain consistent for many reasonable choices of threshold values. We measured accuracy using multiple metrics: (a) mean absolute error (MAE), (b) root mean squared error (RMSE) computed between the predicted and actual water levels, (c) number of time points where the upper or lower thresholds are breached, and (d) the area between water level curves and threshold bars. Table 2 demonstrates that our model GTN outperforms other models with predictions (in red) most closely aligned with the ground truth (in blue) while achieving the lowest MAE and RMSE (in orange). Therefore, we choose our GTN model as Evaluator while training Manager in FIDLAr. In the Appendix, we provide more results in Table 8.
@@
Table 2: Comparison of model performances for the Flood Evaluator on the test set, specifically at time t+1 for measurement station S1. The terms “Over Timesteps” and “Under Timesteps” indicate the number of time steps during which water levels exceed the upper threshold or fall below the lower threshold, respectively. Similarly, “Over Area” and “Under Area” pertain to the area between the water level curve and upper or lower threshold, as was illustrated in Figure 4. Results in orange are the lowest in that column while results in red are the closest to the ground truth (in blue).
Flood Mitigation with FIDLAr.
FIDLAr requires both Evaluator and Manager components. For the Manager model, we experimented with one rule-based method, and two genetic algorithms – one with a physics-based HEC-RAS evaluator (Leon et al. 2020) and one with our DL-based GTN evaluator, and several DL-based managers using MLP, RNN, CNN, GNN, TCN, RCNN, Transformer, and GTN. FIDLAr was measured using (a) the number of time steps where the upper/lower thresholds are exceeded for the water levels, and (b) the area between the water level curves and the threshold bars. Table 3 shows that all DL-based methods consistently performed better for site S1 than rule-based and GA-based approaches. Furthermore, GTN has the best performance under all four metrics, whether it is to control floods or water wastage. The results for all other measurement sites follow a similar pattern, as shown in Table 9 in Appendix C.2.
@@
GA-based Genetic Algorithm - - - -
Genetic Algorithm 86 16.54 454 104
@@
Table 3: Comparison of model performances for the Flood Manager on the test set, specifically at time t+1 for measurement station S1. The denotes that the GA method was used with a physics-based (HEC-RAS) evaluator. The -- denotes that the experiments were timed out. The denotes the GA method was used with the GTN as the evaluator. All other rows are DL-based flood managers with a DL-based GTN as the evaluator. Results in bold are the best in that column.
We visualize water levels for a short period spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for the S1 location. Figure 7 indicates that FIDLAr equipped with GTN model (purple curve) has led to water levels within the upper and lower thresholds, satisfying pre-defined requirements. Moreover, FIDLAr presents the best control (i.e., the least water levels beyond thresholds) for flood mitigation and water waste compared to other baselines. We zoomed in on a 2.5-hour period of the resulting decreased water levels. Its quantitative performance is in Table 10 while more visuals are in Appendix D.
Figure 7: Visualization of water levels with various methods for flood mitigation. We zoomed in $t=2\sim 4.5$ in gray. ‘A’–‘B’ in legend represents the Manager and Evaluator.
@@ 6.3 Ablation Study
The ablation study in Table 4 quantifies the contribution of each component of GTN, by measuring the performance of GTN after removing each of the individual components.
6.4 Analysis of Computational Time
Since FIDLAr was designed for real-time flood control, we measured the running times of the models used in this work. Table 5 shows the running times for the whole flood prediction and mitigation system in its training and test phase. All the DL-based approaches in the test phase are several orders of magnitude faster than the currently used physics-based and GA-based approaches for the flood mitigation task. Rapid inference is a critical property of data-driven DL methods. The table also shows the training times for the DL-based approaches, although they are not necessary for the deployment in reality.
6.5 Explainability
Attention-based methods allow us to calculate the “attention scores” assigned to an input variable to compute a specific output variable. This is exemplified in the heatmap shown in Figure 8, which shows the attention scores assigned to the tide (columns) to compute the gate schedule output (rows) for 24 hours into the future. Note that there are 96 columns and 24 rows because we use 72 hours of past tidal observations and 24 hours of future predicted tidal data to predict 24 hours of the gate schedule into the future. The rows $[0,23]$ correspond to the 24 hours into the future, while the columns $[0,95]$ also include 72 hours of the recent past and 24 hours of the future predicted tidal data. Therefore, $t=72$ corresponds to the “current” time point, and the columns $[72,95]$ correspond to the same time points as rows $[0,23]$.
Figure 8: Importance scores of tide input. $x$ and $y$ axes are the tide and gate over time.
Over Over Under Under
@@ 6.5 Explainability
w/o CNN 37 4.37 476 85.54
w/o Transformer 32 3.57 325 57.42
w/o GNN 56 5.90 479 86.22
w/o LSTM 35 4.34 329 56.74
w/o Attention 32 3.59 341 60.48
GTN (Ours) 22 2.23 299 53.34
Table 4: Ablation study for flood mitigation for the entire test set (for time point t+1 at S1). The last row indicates the performance of the FIDLAr system with GTN as proposed in Figure 6. The best results in the last row are in bold.
Model Flood Prediction Flood Mitigation
@@
GTN (Ours) 119 min 2.95 s 35 min 4.90 s
Table 5: Running time for flood prediction and mitigation. The running time for the rule-based method is not reported since historical data was directly used. GA, which combines a GA-based tool and HEC-RAS, took too long and was not reported, although the time for a small data set is reported in the appendix of our arXiv version. GA, which combines the GA-based tool with GTN, also took too long but was estimated using a smaller sample.
@@ 7 Discussion
The explainability feature, which is shown with an example in Figure 8, can provide significant insights into our results. Firstly, we point out that the brightest patches are in the last 24 columns of the heatmap. Thus FIDLAr pays greater attention to the 24 hours of future predicted tidal data than the past 72 hours, highlighting the importance of forecast information to a DL-based approach to flood mitigation. While tides may have a more predictable pattern over time, the contribution of rain on the water levels can also be seen for other time points. A second critical insight is that the brightest attention patches are in columns where the tide is at its highest is critical to the prediction of gate schedules. Additionally, the water level at the first high tide peak after the “current” time is more significant than the other two. Third, the gate schedule has peaks at times $t=11$ hour and $t=22$ hour into the future, which correspond to the lowest points of the tide. This implies that the optimal time for pre-releasing water is during low tide phases. Opening gates during high tide periods in coastal river systems is less advisable, as it may lead to water flowing back upstream from the ocean. Finally, we observe that there is a light patch around column $t=65$ hour, suggesting mild attention for the previous high tide peak, but almost no attention to any of the peaks before that. This again suggests that we could have chosen to use a smaller window for the past input. Doing this analysis could provide evidence for the right value of $w$, the size of the look-back window.
@@ 8 Conclusions
In this work, we present the shortcomings of the current approaches for flood mitigation and propose FIDLAr, a DL-based tool to address the problem. FIDLAr can compute water “pre-release” schedules for hydraulic structures in a river system to achieve effective and efficient flood mitigation, while ensuring that water wastage is avoided. This was made possible by the use of well-crafted loss functions for the DL models. The dual component design (with a Manager and an Evaluator) is a strength of FIDLAr. It exploits the gradient-based planning and the differentiability of the trained Evaluator model for better optimization. During training, the gradient-based back-propagation from the Evaluator helps to reinforce the Manager.
All the DL-based versions of FIDLAr are several orders of magnitude faster than the (physics-based or GA-based) competitors while achieving improvement over other methods in flood mitigation. These characteristics allow us to entertain the possibility of real-time flood management, which was challenging for previous approaches.
Broader Impact
This research focuses on flood mitigation which has an essential influence on disaster management. As an AI application for social good, our model provides detailed responses on the management of hydraulic structures so as to avoid or mitigate flood events while ensuring minimal water wastage. Our model can also serve as a warning tool. It helps local governments, emergency agencies, and residents to plan well in advance to minimize the losses and dangers caused by impending floods.
@@ Acknowledgments
This work is part of the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) project, which is funded by the National Science Foundation under award number 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views of funding sources.
@@ B.1 Architecture
  • GNN: We consider each time series as one node and the connection as the edge. For the variables in Figure 5, we assign “1” to the edge between two connected nodes, otherwise, “0”. For the architecture, we stacked two graph convolutional networks (GCNs) with $32$ and $16$ neurons. In parallel, a LSTM layer with $32$ units is used to learn the temporal dynamics from the input. We then concatenate the embedding from GCN and LSTM and output the outcomes.
@@
  • Transformer: The standard Transformer encoder (Vaswani et al. 2017) is used. In our case, we only used 1 Encoder with 1 head. We set head size=128, embedding dimension=64, Dense layer with 32 neurons. Dropout is used after each layer above.
  • GTN: We combine the components of GNNs, LSTMs, Transformer, CNNs, and attention, as described in Figure 6. We have $32$ and $16$ for GNN’s channels, $16$ for LSTM units, $96$ for CNN filters, and one Transformer encoder with $3$ heads (head size=192, embedding dimension=96).
@@
When we used the above models for Flood Evaluator and Flood Manager, the input and output need to be changed accordingly (see Figures 2 and 3). To align with real-world scenarios, we enforce stringent constraints that restrict the output of Manager (values for gate and pump openings) to fall within the range of $[MIN,MAX]$ by a modified RELU activation function in Eq. (8). Figure 9 provides the corresponding visualization.
@@ B.2 Training hyper-parameters
Table 6: Best hyperparameters of each DL model for Flood Evaluator.
@@
Table 7: Best hyperparameters of each DL model for Flood Manager.
@@ C.1 Flood prediction at all locations
In Table 2, we provided the comparison of the performance of models for the Flood Evaluator at time point $t+1$, but at one location (S1). In Table 8, we provide the comparison of the performance of different models for the Flood Evaluator on the test set at time $t+1$, but for all locations. Our GTN model still provides the best performance in flood prediction.
Table 8: Comparison of different models for the Flood Evaluator on the test set (at time t+1 for all stations). “Over Timesteps” (and “Under Timesteps”) represent the number of time steps during which the water levels exceed the upper threshold (subceed the lower threshold, resp.). Similarly, “Over Area” (and “Under Area”) refer to the area between the water level curve and the upper threshold (lower threshold, resp.), as shown in Figure 4 in the manuscript.